For a monopolist firm the demand and the total cost functions are given as Q = 20-

0.5P and TC= 4Q2-8Q+15, respectively. 


1
Expert's answer
2021-06-14T13:51:58-0400

Optimum quantity

"Q = 20 - 0.5p"

"0.5p = 20 - Q"

"P = 20-Q \/0.5"

"P = 40 - 2Q"

"TR=P\u00d7Q"

"TR=(40\u22122Q)Q"

"MR = DTR\/DQ"

"MR=40\u22124Q"

"TC=4Q^2 - 8Q"

"MC = DTC\/DQ = 8Q - 8"

Based on equilibrium,

"MR = MC"

"40 - 4Q = 8Q-8"

"40+8= 8Q + 4Q"

"48 = 12Q"

"Q = 48\/12 = 4"

Optimum quantity

"Q = 4"

"P = 40 - 2(Q)"

"P = 40-2(4)"

"P = 40 - 8 = 32"

"TR = P \u00d7 Q = 32 \u00d7 4 = 128"

"TC=4Q^2\n \u22128Q+15"

"TC = 4(4^2) -8(4) + 15"

"TC = 64 - 32 + 15"

"TC = 47"

"Profit = TR-TC"

"Profit = 128-47"

"Profit = 41"